Carraway Seed Company Inc. has for many years cultivated and sold what are known as heritage plants and seeds. For​ example, the company has sought out older varieties of tomato plants that are no longer grown by commercial vegetable farmers since they either take too long to​ mature, do not ship​ well, or do not hold up for long on the store shelf. The company has recently been considering ways to reduce its investment in working capital in order to make itself more profitable. At present the firm has an inventory conversion period of 93 days and the majority of its customers take advantage of its credit terms of 20 days. The company purchases its inventory items on credit terms that allow them 53 days to pay but has always followed a policy of making cash payments for invoices as soon as they are​ received, so the accounts payable deferral period is typically only 13 days.

a.  What are​ Carraway's operating and cash conversion​ cycles?

b.  If Carraway were to decide to take full advantage of its credit terms and delay payment until the last possible​ date, how would this impact their cash conversion​ cycle?

c.  What would be your recommendation to the company with regard to its working capital management practices and​ why?
Answer #1

Answer (a):

Operating cycle = Days Inventory Outstanding (DIO) + Days Sales Outstanding (DSO)

= 93 + 20

= 113 Days

Cash Conversion cycle = Days Inventory Outstanding (DIO) + Days Sales Outstanding (DSO) – Days Payable Outstanding (DPO)

= 93 + 20 - 13

= 100 days

Hence:

Operating cycle = 113 days

Cash Conversion cycle = 100 days

Answer (b):

If Carraway were to decide to take full advantage of its credit terms and delay payment until the last possible​ date, then

Cash Conversion cycle will be = 93 + 20 - 53 = 60 days

Hence:

Impact on Cash Conversion cycle = 100 - 60 = 40 days

If Carraway were to decide to take full advantage of its credit terms and delay payment until the last possible​ date, then Cash Conversion cycle will reduce by 40 days.

Answer (c):

Recommendation to the company with regard to its working capital management practices:

It is stated that the company has recently been considering ways to reduce its investment in working capital in order to make itself more profitable. Hence it should either:

1. Take full advantage of credit period of 53 days available to it and pay its vendors on due date only. This will help reduce its investment in working capital.

OR

2. Talk to its vendors and get suitable cash discount for early payment. It the vendors agree on cash discount it should do cost benefit analysis on whether it will be beneficial to take advantage of cash discount or to pay on due dates and act to implement whichever is beneficial.  This will help it manage its working capital better and improve its profitability.